which term of the arithmetic sequence is -18,given that a1=7 and a2=2?

Answer by chen.aavaz(62)

This is an arithmetic sequence with common difference d=a2-a1=2-7=-5
The nth term is: xn = a + d(n-1) where a is the first term.
Therefore: -18=7-5*(n-1), so n=6